OpenStudy (anonymous):

if one of the roots of the quadratic equation x2+6x+p=0 is twice the other root, find the value of p

OpenStudy (hoblos):

the roots are a &2a then (x-a)(x-2a)=0 x2 -3ax +2a^2 =0 by comparing with the given equation -3a=6 then a=-2 p=2a^2 = 8

OpenStudy (anonymous):

x2+6x+p=0 has r = c , 2c c^2 +6c +p = 0 4c^2 + 12c + p = 0 => 3c^2 + 6c = 0 => 3c (c +2) = 0 => c = 0 and c = -2 c = -2 => 4 -12 = -p Thus p = 8